OpenStudy (anonymous):

limx->1+ (sqrt{x^{2}-x}/x-x^{2})

OpenStudy (anonymous):

\[\lim_{x \rightarrow 1} (\sqrt{x ^{2}-x} ) /x-x ^{2}\]

hartnn (hartnn):

is the question, \(\large \lim \limits_{x \rightarrow 1^+} \dfrac{(\sqrt{x ^{2}-x} )}{ x-x ^{2}} \)

OpenStudy (anonymous):

@hartnn, I would think so, since most of the time homework problems give indeterminate form.

OpenStudy (anonymous):

@hartnn yes it is.

hartnn (hartnn):

ok, then use \(y= \sqrt y \times \sqrt y\) what about \((x-x^2)=...?\) anything cancels out ?

hartnn (hartnn):

or rather, factor out - *minus* first, -(x^2-x)

OpenStudy (anonymous):

i got 1/0.

hartnn (hartnn):

oh, you actually get, \(\large \lim \limits_{x \rightarrow 1^+} \dfrac{1}{-(\sqrt{ x^2-x })}\) right ? now what about the denominator ? when x->1+, x is very very near to 1 but \(\large x>1\) so, denominator will be positive or negative ?

OpenStudy (anonymous):

negative. so is it -infty ?

hartnn (hartnn):

YES, correct :)

OpenStudy (anonymous):

thank you :). i have one more question. can i mention you when i write ?

hartnn (hartnn):

yes, you can, and i saw that limit , tends to 3, right....

OpenStudy (anonymous):

yes. and i have one more which is i'm going to write now.

hartnn (hartnn):

ok, sure.